### v2

@@ 4 Numerical examples
We assess the performance of the sliced sampled completion and the completion based tensor ROM numerically in two examples of parametric dynamical systems. To address the component-wise completion problems for $\boldsymbol{\Phi}^{q}_{\mathbf{i}}$, we use the stable ALS based rank-adaptive tensor completion method in the TT format from grasedyck2019stable , which ensures (37) with prescribed $\varepsilon_{q}$. The corresponding compression $D$-ranks are determined adaptively in the course of completion.
@@ 4.1 Parameterized heat equation
where $u$ is the true (numerical) solution of (59)–(62), while $u_{\ell}$ is the CTROM solution. In order to evaluate CTROM accuracy over the whole parameter domain $\mathcal{A}$ we choose a testing set =${\mathcal{B}}=\{\widetilde{\boldsymbol{\alpha}}_{k}\}_{k=1}^{\widetilde{K}}\subset\mathcal{A}$ and compute the following maximum and mean quantities
@@
${E}_{\text{mean}}$ $1.329\cdot 10^{-3}$ $1.331\cdot 10^{-3}$ $1.338\cdot 10^{-3}$ $4.512\cdot 10^{-4}$ $2.484\cdot 10^{-4}$
@@
We perform five numerical experiments, numbered $1-5$ with parameters and results reported in Table 1. The following quantities were used in all experiments: $N=100$ time-domain snapshots were computed; threshold value $\varepsilon=\sqrt{2}\cdot 10^{-6}$ was set in Algorithm 2 so that the corresponding threshold in Algorithm 1 takes the value of $\varepsilon_{C}=10^{-6}$. The rank-adaptive componenent-wise TT completion for $\boldsymbol{\Phi}^{q}_{i}$, $\mathbf{i}\in\Omega_{C}^{q}$ was ran to ensure (37) with $\varepsilon_{q}=10^{-6}$.
For each experiment we report the $C$-ranks $[q_{1},q_{2}]$ from step 2(c) of Algorithm 1 as well as the following quantities from the the last iteration of Algorithm 2: the sampling rate $|\widetilde{\Omega}_{D}|/|\Omega_{D}|$, the maximum and mean $D$-ranks
@@
In experiments 1-3 the sampling set $\widehat{\mathcal{A}}$ was fixed with $K_{1}=8$, $K_{2}=5$, $K_{3}=5$ and $K_{4}=5$ grid nodes in each parameter direction, respectively. At the same time, the spatial FEM discretization mesh was refined with the maximum element size decreasing from $h=0.33$ to $h=0.2$ and then to $h=0.1$. We observe in the corresponding columns of Table 1 that CTROM is robust with respect to the mesh refinement in all of the reported quantities. This includes the $C$-ranks $[q_{1},q_{2}]$, the maximum and mean $D$-ranks (66), as well as both error quantities (64)–(65) that barely change during mesh refinement.
In experiments 2, 4, 5 the FEM mesh is kept the same with $h=0.2$, but the set $\widehat{\mathcal{A}}$ is refined from $10008\times 5^{3}=$ 1000 to $343010\times 7^{3}=$ 3430 and then to $874812\times 9^{3}=$ 8748 grid nodes, respectively. The purpose of this set of experiments is to show that the sampling rate required for achieving the same level of accuracy decreases as the size of $\widehat{\mathcal{A}}$ grows. Indeed, the sampling rate $|\widetilde{\Omega}_{D}|/|\Omega_{D}|$ first decreases from $0.22$ to $0.094$ and then to $0.048$. At the same time, the $C$-ranks $[q_{1},q_{2}]$ and the maximum and mean $D$-ranks (66) remain stable, while the error quantities (64)–(65) improve (decrease). This demonstrates that slice sampling tensor completion alleviates the curse of dimensionality in that the sampling rate can be decreased for larger $\widehat{\mathcal{A}}$ without the loss of accuracy of the resulting CTROM.
@@ 4.2 Parameterized advection-diffusion equation
We perform the numerical experiments for the advection-diffusion system for varying number of parameters $D$ by setting $\alpha_{j}=0$ for $j=D+1,\ldots,12$. Then, the vectors $[\alpha_{1},\ldots,\alpha_{D}]^{T}$ belong to the parameter domain that is the box $\mathcal{A}=[0.1\pi,0.3\pi]\times[-0.1,0.1]^{D-1}$.
$D$ $6$ $9$ $12$
@@
${E}_{\max}$ 1.8901.890$\cdot$1.890 ⋅ 2.3642.364$\cdot$2.364 ⋅ 1.6441.644$\cdot$1.644 ⋅ 2.2742.274$\cdot$2.274 ⋅ 2.2772.277$\cdot$2.277 ⋅ 1.6531.653$\cdot$1.653 ⋅
@@
${E}_{\text{mean}}$ 1.2421.242$\cdot$1.242 ⋅ 1.0251.025$\cdot$1.025 ⋅ 9.0809.080$\cdot$9.080 ⋅ 9.5339.533$\cdot$9.533 ⋅ 1.4931.493$\cdot$1.493 ⋅ 1.0931.093$\cdot$1.093 ⋅
@@ 5 Conclusions
[…]
@@ Acknowledgements.
M.O. was supported in part by the U.S. National Science Foundation under award DMS-2408978. A.M. and M.O. were supported by the U.S. National Science Foundation under award DMS-2309197. This material is based upon research supported in part by the U.S. Office of Naval Research under award number N00014-21-1-2370 to A.M.

### v3

@@ 4 Numerical examples
We assess the performance of the sliced sampled completion and the completion based tensor ROM numerically in two examples of parametric dynamical systems. To address the component-wise completion problems for $\boldsymbol{\Phi}^{q}_{\mathbf{i}}$, we use SALSA method in the TT format from grasedyck2019stable , which ensures (37) with prescribed $\varepsilon_{q}$. All other input parameters of SALSA are left at default values. The corresponding compression $D$-ranks are determined adaptively during the completion process. We note that SALSA method does not necessarily produce a minimal-rank TT tensor that fits the data within the prescribed tolerance, and thus it can be viewed as an approximate solution to the completion problem (7).
The code of Matlab implementation of CTROM used for the numerical experiments is available upon request.
@@ 4.1 Parameterized heat equation
where $u$ is the true (numerical) solution of (59)–(62), while $u_{\ell}$ is the CTROM solution. In order to evaluate CTROM accuracy over the whole parameter domain $\mathcal{A}$ we choose a testing set ${\mathcal{B}}=\{\widetilde{\boldsymbol{\alpha}}_{k}\}_{k=1}^{\widetilde{K}}\subset\mathcal{A}$ and compute the following maximum and mean quantities
@@
${E}_{\text{mean}}$ $29\cdot 10^{-3} 1.331$$31\cdot 10^{-3}$ $1.338\cdot 10^{-3}$ $4.512\cdot 10^{-4}$ $2.484\cdot 10^{-4}$
@@
We perform five numerical experiments, numbered 151-5 with parameters and results reported in Table 1. The following quantities were used in all experiments: $N=100$ time-domain snapshots were computed; threshold value $\varepsilon=\sqrt{2}\cdot 10^{-6}$ was set in Algorithm 2 so that the corresponding threshold in Algorithm 1 takes the value of $\varepsilon_{C}=10^{-6}$. The rank-adaptive componenent-wise TT completion for $\boldsymbol{\Phi}^{q}_{i}$, $\mathbf{i}\in\Omega_{C}^{q}$ was ran to ensure (37) with $\varepsilon_{q}=10^{-6}$.
For each experiment we report the $C$-ranks $q_{1},q_{2}]$ from step 2(c) of Algorithm 1 as well as the following quantities from the the last iteration of Algorithm 2: the sampling rate $|\widetilde{\Omega}_{D}|/|\Omega_{D}|$, the maximum and mean $D$-ranks
@@
In experiments 1-3 the sampling set $\widehat{\mathcal{A}}$ was fixed with $K_{1}=8$, $K_{2}=5$, $K_{3}=5$ and $K_{4}=5$ grid nodes in each parameter direction, respectively. At the same time, the spatial FEM discretization mesh was refined with the maximum element size decreasing from $h=0.33$ to $h=0.2$ and then to $h=0.1$. We observe in the corresponding columns of Table 1 that CTROM is robust with respect to the mesh refinement in all of the reported quantities. This includes the $C$-ranks $q_{1},q_{2}]$, the maximum and mean $D$-ranks (66), as well as both error quantities (64)–(65) that barely change during mesh refinement.
In experiments 2, 4, 5 the FEM mesh is kept the same with $h=0.2$, but the set $\widehat{\mathcal{A}}$ is refined from $8\times 5^{3}=1000$ to $10\times 7^{3}=3430$ and then to $12\times 9^{3}=8748$ grid nodes, respectively. The purpose of this set of experiments is to show that the sampling rate required for achieving the same level of accuracy decreases as the size of $\widehat{\mathcal{A}}$ grows. Indeed, the sampling rate $|\widetilde{\Omega}_{D}|/|\Omega_{D}|$ first decreases from $0.22$ to $0.094$ and then to $0.048$. At the same time, the $C$-ranks $q_{1},q_{2}]$ and the maximum and mean $D$-ranks (66) remain stable, while the error quantities (64)–(65) improve (decrease). This demonstrates that slice sampling tensor completion alleviates the curse of dimensionality in that the sampling rate can be decreased for larger $\widehat{\mathcal{A}}$ without the loss of accuracy of the resulting CTROM.
In the final experiment for the parameterized heat problem, we investigate
Figure 2: CTROM solution error versus completion accuracy.
the dependence of the ROM solution error on the completion accuracy. The ROM solution error is defined in equations (64)–(65). The completion error is characterized by $\varepsilon$ in (46), and the estimate (38) provides an upper bound for it. Since we do not have direct access to $\varepsilon$ or to certain quantities appearing in (38), the completion accuracy is instead assessed on a test set, as described in (47). The results presented in Fig. 2 show that the ROM errors decrease almost linearly over the given range of $\varepsilon$. In this experiment, we set $h=0.2$, used a $10\times 7^{3}$ mesh in $\widehat{\mathcal{A}}$, and chose $\ell=q_{2}$.
@@ 4.2 Parameterized advection-diffusion equation
We perform the numerical experiments for the advection-diffusion system for varying number of parameters $D$ by setting $\alpha_{j}=0$ for $j=D+1,\ldots,12$. Then, the vectors $\alpha_{1},\ldots,\alpha_{D}]^{T}$ belong to the parameter domain that is the box $\mathcal{A}=[0.1\pi,0.3\pi]\times[-0.1,0.1]^{D-1}$.
$D$ 66 99 $12$
@@
${E}_{\max}$ $1.890\cdot$ $2.364\cdot$ $1.644\cdot$ $4\cdot 2.2772.2$$7\cdot$ $1.653\cdot$
@@
${E}_{\text{mean}}$ $1.242\cdot$ $1.025\cdot$ $9.080\cdot$ $9.533\cdot$ $1.493\cdot$ $1.093\cdot$
@@
The final experiment for the advection-diffusion equation is to compare the performance of CTROM to the conventional TROM that uses the full tensor of snapshots, as introduced in mamonov2022interpolatory . For this comparison the setup is as follows. We use $D=9$ parameters and the grid $^{\widehat{\mathcal{A}}}$ with $K_{1}=5$ and $K_{j}=3$, $j=2,\ldots,D$, for a total of $32,805$ grid points. While $^{\widehat{\mathcal{A}}}$ is rather coarse, it is still the finest grid that can be handled by the conventional TROM for $D=9$ on the machine used for the numerical experiments. We set $\varepsilon=5\cdot 10^{-4}$ in Algorithm 2 and we use the same accuracy for TT decomposition in the offline stage of the conventional TROM. We compare the two TROM aproaches by timing the offline stage of both, since it dominates the computational cost. For the conventional TROM the computation of the snapshots for the tensor $\boldsymbol{\Phi}$ took 2828hrs 5151min with additional 1414min to perform its TT decomposition for a total of 2929hrs 0505min for the offline stage. The offline stage of CTROM, Algorithm 2, took 77hrs 1010min which includes computing $4,172$ samples, i.e., the percentage of observed entries of $\boldsymbol{\Phi}$ is $12.71\%$. Thus, for this particular case we observe more than four-fold gain in computational cost of CTROM compared to the conventional TROM. Note also that the relatively high observed entry percentage in Algorithm 2 is due to a coarse grid $^{\widehat{\mathcal{A}}}$. One should expect even greater performance gains of CTROM for finer grids $^{\widehat{\mathcal{A}}}$.
@@ 5 Conclusions
Applying the completion method to two linear parametric parabolic PDEs, discretized via the finite element method, revealed the following properties: (i) The resulting tensor ranks are insensitive to discretization parameters in space and time, as well as to grid resolution in the parameter domain. (ii) The ranks increase with more accurate completion, requiring additional memory for storing the recovered tensor. This dependency is of the form $\text{d.o.f.}=O(\varepsilon^{-\alpha})$, where $\varepsilon$ represents the completion accuracy and $\alpha\geq 0$ is an exponent. In our tests, $\alpha$ did not exceed $\frac{2}{3}$. (iii) The total number and percentage of observed tensor entries required for successful completion depends on the target accuracy, but more significantly on the dimension of the parameter space and on how fine is the parameter domain grid. Specifically, the total number of entries required increases roughly linearly with the number of nodes for any parameter. Therefore, if the grid in $\mathcal{A}$ is refined by doubling the nodes in each direction, the number of observed entries required doubles, while the percentage of observed entries ‘decrease’ by a factor of $2^{D-1}$. This explains our ability to recover snapshot tensors from fewer than $0.01\%$ of entries in higher-dimensional parameter spaces. (iv) The compression achieved by the HTT format for snapshot tensors similarly depends on parameter space dimension and resolution. In our numerical examples, we achieved a compression factor of $10^{6}$ while preserving an accuracy of $10^{-4}$.
@@ Acknowledgements.
A.M. and M.O. were supported by the U.S. National Science Foundation under award DMS-2309197. This material is based upon research supported in part by the U.S. Office of Naval Research under award number N00014-21-1-2370 to A.M.